# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v3

@@ 8 Conclusions
To ensure the reproducibility of our work, we release all related code, dataset metadata, and detailed experimental configurations. Our evaluation framework AudioTrust, the automated evaluation scripts, and the leaderboard generation code are hosted in an anonymous GitHub repository: https://github.com/AudioTrust/AudioTrust. Appendix C describes the decoupled two-stage architecture (inference and evaluation) of our platform, which provides guidance for reproducing our evaluation pipeline. Throughout the main paper (Sections 2–7), we explicitly indicate the corresponding appendix sections. Specifically, Appendices D through I present detailed treatments of six evaluation dimensions: fairness, hallucination, safety, privacy, robustness, and verification, which cover attack strategies, dataset construction, evaluation protocols, and the metrics employed. For instance, Appendices D.1 and D.2 describe the dataset construction process for fairness evaluation, while Appendices E.3.1 and E.3.2 explain the details of the metrics for hallucination evaluation. We believe that the public release of these resources will facilitate both reproducibility and extension of our study by the community.
@@ Appendix B Benchmark Models
To systematically investigate these trustworthiness aspects, we have selected a diverse set of models. This set includes both mainstream proprietary commercial models, such as GPT-4 (OpenAI et al., 2024) and Gemini (Team et al., 2023), as well as representative and robust open-source ALLMs, including Qwen2-Audio (Chu et al., 2024) and MiniCPM-o 2.6 (Yao et al., 2024). To ensure fairness and objectivity, all models are systematically tested on the same datasets and with identical evaluation metrics, followed by thorough comparative analyses of experimental results. It is worth noting that our methodology considers not only the fundamental audio comprehension capabilities of each model, but also examines their potential strengths and limitations in aspects such as complex interactions and knowledge transfer. This systematic safety evaluation provides a solid foundation for the future optimization and development of ALLMs.
@@ B.1 Open-Source Models
1. SALMONN (Tang et al., 2024) pioneered a dual-encoder architecture (Whisper speech encoder and BEATs audio encoder) together with a window-level Q-Former and LoRA adapters. This enables the pretrained Vicuna text LLM to achieve unified understanding of speech, environmental sounds, and music. The model also demonstrates emergent capabilities in cross-modal reasoning beyond the training tasks and in few-shot activation tuning.
@@
2. Ultravox (AI, 2024) directly maps raw audio into the high-dimensional representation space of LLMs, thereby seamlessly eliminating the traditional ASR stage. This model not only comprehends speech content but also captures paralinguistic features such as tone and pauses, and supports streaming text outputs.
@@
3. Qwen2-Audio (Chu et al., 2024) is a large-scale audio-language model that establishes a seamless pipeline between the Whisper-large-v3 encoder and the Qwen-7B language model, thereby supporting both spoken dialogue and audio analysis interaction modes. In real conversational and multitask zero-shot evaluations, the model leverages Mel-spectrograms of 16kHz audio combined with instruction tuning and Direct Preference Optimization (DPO), significantly improving the precision and robustness of responses to human intent.
@@
4. MiniCPM-o 2.6 (Yao et al., 2024) integrates four major components: SigLip-400M, Whisper-medium, ChatTTS-200M, and Qwen2.5-7B, supporting bilingual real-time dialogue in an end-to-end multimodal fashion, as well as controllable interactions in emotion and speaking rate, and high-quality voice cloning. It consistently outperforms proprietary models of equivalent scale on benchmarks such as OpenCompass and StreamingBench.
@@
5. Step-Fun (Huang et al., 2025) is a production-ready open-source real-time speech–text multimodal system that tackles data-collection cost, weak dynamic control, and limited intelligence via four pillars: a 130B unified understanding–generation model, a generative speech data engine enabling affordable voice cloning and distilling the lightweight Step-Audio-TTS-3B, an instruction-driven fine-control mechanism spanning dialects, emotions, singing, and rap, and an enhanced cognitive layer with tool calling and role-playing for complex tasks.
@@
6. Qwen2.5-Omni (Xu et al., 2025b) builds upon Qwen2.5-VL/Audio by introducing the Thinker-Talker architecture and TMRoPE (Time-aligned Multimodal RoPE) temporal alignment embedding. This allows the model to stream and process text, image, audio, and video inputs concurrently within a single framework, with the ability to produce both textual and natural speech outputs in synchronization.
@@
7. Kimi-Audio (Ding et al., 2025) is an open-source audio foundation model for understanding, generation, and conversation; it adopts a 12.5Hz audio tokenizer and an LLM-based architecture that ingests continuous features and emits discrete tokens, alongside a chunk-wise streaming detokenizer via flow matching for low-latency inference.
@@
8. OpenS2S (Wang et al., 2025a) built on the BLSP-Emo empathetic speech-to-text backbone, it introduces a streaming interleaved decoding architecture for low-latency speech generation while capturing rich paralinguistic cues for expressive responses.
@@
9. Step-Audio 2 (Wu et al., 2025) is an end-to-end multimodal LLM for industry-grade audio understanding and speech conversation, combining a latent audio encoder with reasoning-centric RL to boost ASR and audio comprehension; by folding discrete audio-token generation into language modeling, it becomes highly responsive to paralinguistic cues in real time.
@@ B.2 Closed-Source Models
10. Gemini-1.5 Pro leverages a Mixture-of-Experts architecture for unified reasoning across speech, image, and text. It supports audio inputs up to 19 hours in duration and contexts up to the million-token scale, enabling seamless processing for tasks such as audio summarization, transcription, and translation.
11. GPT-4o Audio is the first developer-oriented interactive audio model that supports both understanding and generation of speech. It is capable of speech transcription, summarization, sentiment analysis, and conversational dialogue.
12. GPT-4o mini Audio is designed to deliver cost-effective yet robust audio understanding and generation. It supports a variety of audio input formats and can produce seamless bimodal (text and speech) output with customizable speech styles, making it applicable to edge devices and large-scale embedded deployments.
13. Gemini-2.5 Flash retains the core multimodal design of the Pro version while significantly optimizing inference speed and computational efficiency. This version supports up to 8.4 hours of audio input and million-token context windows, with dramatically reduced latency and operational cost compared to the Pro variant, while still covering tasks like audio summarization, transcription, and translation.
14. Gemini-2.5 Pro further advances multimodal reasoning, introducing a dynamic “thinking budget” mechanism that adaptively allocates computational resources based on instruction and system constraints. Its superior performance on video understanding benchmarks extends to the audio domain, enabling streaming responses for complex tasks such as conversational QA, scenario retrieval, and reasoning through efficient temporal alignment and cross-modal integration.
@@ D.2 Dataset Construction Method
During the data construction phase, we utilized the multi-modal generation model GPT-4o (OpenAI et al., 2024) to produce textual content, which was then transformed into audio using the F5-TTS model (Chen et al., 2024). The entire data generation process begins with a text-based prompt, constructed entirely by GPT-4o to ensure natural language fluency and semantic consistency. To enable controlled experiments with single-variable variation, each data pair consists of two audio samples with identical semantic content, differing solely in their associated sensitive attributes.
@@ D.3 Evaluation Method and Metrics
We utilized statistical methods and developed the unfairness metrics to quantitatively assess model fairness (Xu et al., 2025a; Deldjoo & Nazary, 2024). Let an audio–language model be , mapping the audio–text input space to the text output space . Denote by the set of sensitive attribute values (e.g. ). For a test collection , we introduce a discriminator detecting sensitive attributes in generations.
Group Unfairness Score. For any group label , define the group unfairness score
@@ Results.
| Gemini-1.5 Pro | 0.65 | 0.10 | 0.75 | 0.10 |
| Gemini-2.5 Flash | 0.80 | 0.50 | 0.89 | 0.80 |
| Gemini-2.5 Pro | 1.00 | 0.50 | 1.00 | 1.00 |
| GPT-4o Audio | 0.80 | 0.70 | 0.75 | 0.80 |
| GPT-4o Mini Audio | 0.70 | 0.90 | 0.80 | 0.50 |
@@
| Gemini-1.5 Pro | 1.00 | 1.00 | 0.00 | 1.00 |
| Gemini-2.5 Flash | 1.00 | 0.80 | 0.80 | 1.00 |
| Gemini-2.5 Pro | 1.00 | 1.00 | 1.00 | 1.00 |
| GPT-4o | 1.00 | 1.00 | 0.80 | 1.00 |
| GPT-4o Mini | 1.00 | 1.00 | 1.00 | 0.80 |
| Average | 0.89 | 0.75 | 0.83 | 0.82 |
@@ E.3.1 Experimental Design
We designed a comprehensive red-teaming framework to evaluate hallucination tendencies in ALLMs across four distinct categories: physical violations, label mismatches, logical violation, and content mismatches. Our methodology involved creating specialized datasets that test specific aspects of model robustness against hallucination. For physical violations, we synthesized sounds that contradict fundamental laws (e.g., underwater flames); for label mismatches, we deliberately paired audio with incorrect attribute descriptions (e.g., mislabeling speaker gender); for temporal disorders, we manipulated cause-and-effect sequences in audio events (e.g., reversing “gunshot followed by falling”); and for content mismatches, we paired audio with semantically contradictory descriptions (e.g., labeling bird chirping as dog barking). We conducted four experimental protocols with varying prompt structures:
@@ E.4 Evaluation Methodology
In the final metrics calculation stage, a series of quantitative metrics are utilized to objectively and thoroughly evaluate model performance (see Section˜E.3.2 for details). Specifically, these metrics include: the HDR, which measures the proportion of true physical violations accurately identified by the model, thus reflecting its detection sensitivity; the FHR, which assesses the proportion of misclassifications the model makes in normal control cases without actual violations, hence indicating the model’s robustness and false positive rate. Additionally, we introduce the EQS, which is assigned by three expert human raters on a 5-point scale. Ratings are given from multiple perspectives, including physical correctness, the relevance of the explanation to the audio facts, and the logic and clarity of the reasoning process. The final EQS is computed as the mean score across all raters and all test cases, thereby providing a comprehensive quantitative measure of the model’s interpretability. Overall, this multi-dimensional evaluation framework effectively captures the model’s Overall competence in the context of physical violation detection, encompassing detection accuracy, false positives, and explanation quality, thus offering a reliable experimental foundation for subsequent optimization and improvement of the methods.

### v4

@@ 8 Conclusions
Acknowledgments
This research was supported by the National Research Foundation, Singapore, and the Cyber Security Agency of Singapore under the National Cybersecurity R&D Programme and the CyberSG R&D Programme Office (Award CRPO-GC3-NTU-001), NTU-NAP startup grant (024584-00001), NTU startup funding (025559-00001), and the Singapore Ministry of Education Tier 1 Grant (RG19/25). The authors from Tsinghua University are partially supported by the National Natural Science Foundation of China (No. U2341228).
Any opinions, findings, conclusions, or recommendations expressed in these materials are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore, the Cyber Security Agency of Singapore, or the CyberSG R&D Programme Office.
@@ Acknowledgments
To ensure the reproducibility of our work, we release all related code, dataset metadata, and detailed experimental configurations. Our evaluation framework AudioTrust, the automated evaluation scripts, and the leaderboard generation code are hosted in an GitHub repository: https://github.com/JusperLee/AudioTrust. Appendix C describes the decoupled two-stage architecture (inference and evaluation) of our platform, which provides guidance for reproducing our evaluation pipeline. Throughout the main paper (Sections 2–7), we explicitly indicate the corresponding appendix sections. Specifically, Appendices D through I present detailed treatments of six evaluation dimensions: fairness, hallucination, safety, privacy, robustness, and verification, which cover attack strategies, dataset construction, evaluation protocols, and the metrics employed. For instance, Appendices D.1 and D.2 describe the dataset construction process for fairness evaluation, while Appendices E.3.1 and E.3.2 explain the details of the metrics for hallucination evaluation. We believe that the public release of these resources will facilitate both reproducibility and extension of our study by the community.
@@ Appendix B Benchmark Models
To systematically investigate these trustworthiness aspects, we have selected a diverse set of models. This set includes both mainstream proprietary commercial models, such as GPT-4 (OpenAI et al., 2024) and Gemini (Team et al., 2023), as well as representative and robust open-source ALLMs, including Qwen2-Audio (Chu et al., 2024) and MiniCPM-o 2.6 (Yao et al., 2024). To ensure fairness and objectivity, all models are systematically tested on the same datasets and with identical evaluation metrics, followed by thorough comparative analyses of experimental results. It is worth noting that our methodology considers not only the fundamental audio comprehension capabilities of each model, but also examines their potential strengths and limitations in aspects such as complex interactions and knowledge transfer. This systematic safety evaluation provides a solid foundation for the future optimization and development of ALLMs. To vividly visualize these comparative strengths and weaknesses, we present a six-dimensional trustworthiness profile for each model (see Figure 3), covering fairness (F), hallucination (H), safety (S), privacy (P), robustness (R), and authentication (A). In the following descriptions, we analyze the specific radar chart performance of each model to highlight their distinct trade-offs and alignment characteristics.
@@ B.1 Open-Source Models
1. SALMONN (Tang et al., 2024) pioneered a dual-encoder architecture (Whisper speech encoder and BEATs audio encoder) together with a window-level Q-Former and LoRA adapters. This enables the pretrained Vicuna text LLM to achieve unified understanding of speech, environmental sounds, and music. The model also demonstrates emergent capabilities in cross-modal reasoning beyond the training tasks and in few-shot activation tuning. The radar chart reveals a notable contradiction in the design of SALMONN: although it achieves surprisingly high scores in privacy protection as an open-source model, this comes at a significant cost to fairness and reliability. We hypothesize that SALMONN’s relatively weak capability in environmental sound perception limits its performance in privacy inference tasks, which in turn leads to artificially elevated privacy scores.
@@
2. Ultravox (AI, 2024) directly maps raw audio into the high-dimensional representation space of LLMs, thereby seamlessly eliminating the traditional ASR stage. This model not only comprehends speech content but also captures paralinguistic features such as tone and pauses, and supports streaming text outputs. Ultravox demonstrates excellent safety and decent hallucination resistance, but its robustness is critically low. This indicates that while the model is well-aligned semantically, its end-to-end architecture is extremely fragile to acoustic perturbations.
@@
3. Qwen2-Audio (Chu et al., 2024) is a large-scale audio-language model that establishes a seamless pipeline between the Whisper-large-v3 encoder and the Qwen-7B language model, thereby supporting both spoken dialogue and audio analysis interaction modes. In real conversational and multitask zero-shot evaluations, the model leverages Mel-spectrograms of 16kHz audio combined with instruction tuning and Direct Preference Optimization (DPO), significantly improving the precision and robustness of responses to human intent. The cross-dimensional analysis highlights a significant discrepancy in Qwen2-Audio: although it performs strongly in terms of safety and authentication, its privacy score remains alarmingly low. This indicates that the model can effectively identify speakers but fails to safeguard their sensitive attributes. A potential reason may lie in the lack of effective safety-alignment training after integrating the new modality, or an excessive reliance on ASR outputs, which leads to insufficient protection of sensitive audio attributes.
@@
4. MiniCPM-o 2.6 (Yao et al., 2024) integrates four major components: SigLip-400M, Whisper-medium, ChatTTS-200M, and Qwen2.5-7B, supporting bilingual real-time dialogue in an end-to-end multimodal fashion, as well as controllable interactions in emotion and speaking rate, and high-quality voice cloning. It consistently outperforms proprietary models of equivalent scale on benchmarks such as OpenCompass and StreamingBench. The model exhibits a clear trade-off: it prioritizes robustness and safety while largely neglecting privacy and performing poorly in controlling hallucinations. This characteristic suggests that the system demonstrates strong resistance to noise and harmful prompts yet is prone to private information leakage and the generation of unsupported content.
@@
5. Step-Fun (Huang et al., 2025) is a production-ready open-source real-time speech–text multimodal system that tackles data-collection cost, weak dynamic control, and limited intelligence via four pillars: a 130B unified understanding–generation model, a generative speech data engine enabling affordable voice cloning and distilling the lightweight Step-Audio-TTS-3B, an instruction-driven fine-control mechanism spanning dialects, emotions, singing, and rap, and an enhanced cognitive layer with tool calling and role-playing for complex tasks. Step-Fun stands out as the most balanced open-source model, particularly in terms of fairness and robustness. Unlike comparable models that often sacrifice fairness for safety, Step-Fun maintains moderate performance across all dimensions, although its authentication results indicate remaining room for improvement in identity verification.
@@
6. Qwen2.5-Omni (Xu et al., 2025b) builds upon Qwen2.5-VL/Audio by introducing the Thinker-Talker architecture and TMRoPE (Time-aligned Multimodal RoPE) temporal alignment embedding. This allows the model to stream and process text, image, audio, and video inputs concurrently within a single framework, with the ability to produce both textual and natural speech outputs in synchronization. Qwen2.5-Omni enhances safety but exacerbates cross-dimensional imbalance: fairness decreases while privacy becomes nearly negligible. This reflects a broader trend wherein cross-modal alignment training may inadvertently suppress the fine-grained reasoning capabilities necessary for achieving both fairness and privacy protection.
@@
7. Kimi-Audio (Ding et al., 2025) is an open-source audio foundation model for understanding, generation, and conversation; it adopts a 12.5Hz audio tokenizer and an LLM-based architecture that ingests continuous features and emits discrete tokens, alongside a chunk-wise streaming detokenizer via flow matching for low-latency inference. Kimi-Audio exhibits the most extreme alignment trade-off in our benchmark. While it achieves near-perfect safety, such over-cautious behavior likely results in the lowest fairness score and very limited privacy protection. This may occur because the model indiscriminately rejects legitimate queries from marginalized groups or fails to distinguish between safe and private contexts.
@@
8. OpenS2S (Wang et al., 2025a) built on the BLSP-Emo empathetic speech-to-text backbone, it introduces a streaming interleaved decoding architecture for low-latency speech generation while capturing rich paralinguistic cues for expressive responses. Interestingly, OpenS2S exhibits characteristics opposite to most models: it performs strongly in hallucination resistance and authentication but shows weaker robustness and fairness. This suggests that the model achieves high precision under clean conditions yet lacks the generalization ability to handle diverse speakers or noisy environments.
@@
9. Step-Audio 2 (Wu et al., 2025) is an end-to-end multimodal LLM for industry-grade audio understanding and speech conversation, combining a latent audio encoder with reasoning-centric RL to boost ASR and audio comprehension; by folding discrete audio-token generation into language modeling, it becomes highly responsive to paralinguistic cues in real time. The model performs reasonably well in robustness and safety but exhibits systematic weaknesses in fairness and privacy. This indicates that its reasoning-centered reinforcement learning optimization primarily focuses on utility and ASR performance rather than ethical alignment.
@@ B.2 Closed-Source Models
10. Gemini-1.5 Pro leverages a Mixture-of-Experts architecture for unified reasoning across speech, image, and text. It supports audio inputs up to 19 hours in duration and contexts up to the million-token scale, enabling seamless processing for tasks such as audio summarization, transcription, and translation. Although Gemini-1.5 Pro demonstrates the best safety performance and excels in robustness and authentication, its relatively low hallucination score reveals a critical vulnerability. The model exhibits strong resistance to adversarial attacks but tends to fabricate content, highlighting a notable gap between safety alignment and factual grounding.
11. GPT-4o Audio is the first developer-oriented interactive audio model that supports both understanding and generation of speech. It is capable of speech transcription, summarization, sentiment analysis, and conversational dialogue. GPT-4o Audio demonstrates the most comprehensive performance, leading in authentication and achieving high scores in safety, privacy, and fairness. However, similar to Gemini, it remains affected by hallucinations, suggesting that scaling up the model can enhance nearly all dimensions of trustworthiness, while hallucination issues may stem from inaccuracies in the training data.
12. GPT-4o mini Audio is designed to deliver cost-effective yet robust audio understanding and generation. It supports a variety of audio input formats and can produce seamless bimodal (text and speech) output with customizable speech styles, making it applicable to edge devices and large-scale embedded deployments. The "mini" variant illustrates a clear compression trade-off: it retains the superior authentication and safety of the larger model (even improves privacy), but its robustness collapses to 0.228. This suggests that lightweight models can be safe and private but lack the parameter redundancy to handle adversarial noise.
13. Gemini-2.5 Flash retains the core multimodal design of the Pro version while significantly optimizing inference speed and computational efficiency. This version supports up to 8.4 hours of audio input and million-token context windows, with dramatically reduced latency and operational cost compared to the Pro variant, while still covering tasks like audio summarization, transcription, and translation. Gemini-2.5 Flash prioritizes safety and robustness to serve as a reliable production model. However, cross-dimensional analysis shows it lags in hallucination and fairness, suggesting that optimization for speed and safety may have compromised its ability to handle nuanced equity tasks and factual verification.
14. Gemini-2.5 Pro further advances multimodal reasoning, introducing a dynamic “thinking budget” mechanism that adaptively allocates computational resources based on instruction and system constraints. Its superior performance on video understanding benchmarks extends to the audio domain, enabling streaming responses for complex tasks such as conversational QA, scenario retrieval, and reasoning through efficient temporal alignment and cross-modal integration. Compared to its predecessor, Gemini-2.5 Pro has improved hallucination resistance while maintaining elite safety and robustness. Nevertheless, fairness remains a persistent challenge, reinforcing the observation that advanced reasoning capabilities do not automatically translate to equitable decision-making without explicit intervention.
@@ D.2 Dataset Construction Method
During the data construction phase, we utilized the multi-modal generation model GPT-4o (OpenAI et al., 2024) to produce textual content, which was then transformed into audio using the F5-TTS model (Chen et al., 2025). The entire data generation process begins with a text-based prompt, constructed entirely by GPT-4o to support natural language fluency and semantic consistency. To enable controlled experiments with single-variable variation, each data pair consists of two audio samples with identical semantic content, differing solely in their associated sensitive attributes.
@@ D.3 Evaluation Method and Metrics
We utilized statistical methods and developed the fairness metrics to quantitatively assess model fairness (Xu et al., 2025a; Deldjoo & Nazary, 2024). Let an audio large language model be , mapping the audio–text input space to the text output space . Denote by the set of sensitive attribute values (e.g. ). For a test collection , we introduce a discriminator detecting sensitive attributes in generations.
Group Fairness Score. For any group label , define the group fairness score
@@ Results.
| Gemini-1.5 Pro | 0.35 | 0.90 | 0.25 | 0.90 |
| Gemini-2.5 Flash | 0.20 | 0.50 | 0.11 | 0.20 |
| Gemini-2.5 Pro | 0.00 | 0.50 | 0.00 | 0.00 |
| GPT-4o Audio | 0.20 | 0.30 | 0.25 | 0.20 |
| GPT-4o Mini Audio | 0.30 | 0.10 | 0.20 | 0.50 |
@@
| Gemini-1.5 Pro | 0.00 | 0.00 | 1.00 | 0.00 |
| Gemini-2.5 Flash | 0.00 | 0.20 | 0.20 | 0.00 |
| Gemini-2.5 Pro | 0.00 | 0.00 | 0.00 | 0.00 |
| GPT-4o | 0.00 | 0.00 | 0.20 | 0.00 |
| GPT-4o Mini | 0.00 | 0.00 | 0.00 | 0.20 |
| Average | 0.11 | 0.25 | 0.17 | 0.18 |
@@ E.3.1 Experimental Design
We designed a comprehensive red-teaming framework to evaluate hallucination tendencies in ALLMs across four distinct categories: physical violations, label mismatches, logical violation, and content mismatches. Our methodology involved creating specialized datasets that test specific aspects of model robustness against hallucination. For physical violations, we synthesized sounds that contradict fundamental laws (e.g., underwater flames); for label mismatches, we deliberately paired audio with incorrect attribute descriptions (e.g., mislabeling speaker gender); for temporal disorders, we manipulated cause-and-effect sequences in audio events (e.g., reversing “gunshot followed by falling”); and for content mismatches, we paired audio with semantically contradictory descriptions (e.g., labeling bird chirping as dog barking). At the same time, we found that if the model’s response relies on overly forced explanations (for example, "The object landed precisely on a perfect sound-absorbing material"), to avoid ambiguity and excessive rationalization, this evaluation does not rely on the open-ended responses of the model. We use GPT-4o as the automatic evaluator, which follows pre-defined common sense rules. This evaluator is explicitly instructed to ignore special environments or recording settings and only focus on whether the described event follows the laws of everyday physics. Models that unquestioningly accept extreme or temporary set scenarios will be regarded as having produced hallucinations and will receive a lower score. We conducted four experimental protocols with varying prompt structures:
@@ E.4 Evaluation Methodology
In the metrics calculation stage, we utilize a series of quantitative metrics to evaluate model performance (see Section˜E.3.2 for complete formulations). Our evaluation framework is twofold, comprising both automated metrics and human-based judgments.First, we programmatically compute two metrics by parsing the model’s textual output and comparing it against ground-truth labels, requiring no human judgment. These are: 1) The HDR (Hallucination Detection Rate), which measures detection sensitivity by calculating the proportion of true physical violations (our positive class) that the model correctly identifies. 2) The FHR (False Hallucination Rate), which assesses the false positive rate by calculating the proportion of normal control cases (our negative class) that the model incorrectly flags as violations.Second, to complement these automated metrics, we introduce the EQS (Explanation Quality Score) as a human-based measure of interpretability. This score is assigned by three expert human raters on a 5-point scale. Ratings evaluate multiple perspectives, including physical correctness, the relevance of the explanation to the audio facts, and the logic and clarity of the reasoning process. The final EQS is computed as the mean score across all raters and test cases.
Overall, this multi-dimensional framework, combining automated detection accuracy (HDR/FHR) with human-rated explanation quality (EQS), effectively captures the model’s competence and provides a reliable experimental foundation.